Three recent graduates of the computer science program at the University of Tennessee are forming a company that will write and distribute new application software for the iPhone. ​ Initially, the corporation will operate in the southern region of​ Tennessee, Georgia, North​ Carolina, and South Carolina. A small group of private investors in the​ Atlanta, Georgia area is interested in financing the startup company and two financing plans have been put forth for​ consideration:

** The first (Plan A) is an all-common-equity capital structure. $2.0million dollars would be raised by selling common stock at $20.00 per common share.

**Plan B would involve the use of financial leverage. $1.0 million dollars would be raised by selling bonds with an effective interest rate of 11.0% (per annum), and the remaining $1.0 million would be raised by selling common stock at the $20.00 price per share. The use of financial leverage is considered to be a permanent part of the​ firm's capitalization, so no fixed maturity date is needed for the analysis. A 30% tax rate is deemed appropriate for the analysis.

A - The EBIT indifference level associated with the two financing plans is?

B - Using EBIT of $300,000, complete the segment of the income statement for Plan A​ below:  ​(Round income statement amounts to the nearest dollar except the EPS to the nearest​ cent.)

EBIT
Less: Interest Expense
Earnings Before Taxes
Less: Taxes @ 30%
Net Income
Number of Common Shares
EPS
Answer #1

Plan A. No.of shares XA = $ 2 million/ $20 = 100,000

Plan B : No.of shares XB = $ 1 million/ $20 = 50,000

EBIT indifference point between plan A and B is given by the formula

(EBIT - IA) * (1-t) / XA =  (EBIT - IB) * (1-t) / XB

where IA and IB are the interest expenses in plan A and plan B

IA = 0 (as there is no debt)

IB = 11% of $ 1 million = $110,000

So, (EBIT -0 )* (1-0.3) / 100000 = ( EBIT - 110000)* (1-0.3) / 50000

=> EBIT = 2 * (EBIT- 110000)

=> EBIT = $220,000

This is the EBIT indifference level associated with the two financing plans

2. For Plan A (all equity)

EBIT 300,000

Less: interest expense 0

Earnings before taxes 300,000

Less: Taxes @30% 90,000

Net Income 210,000

No. of Common shares 100,000

EPS 2.10